A resistor of 50 ohm on inductor of 20/π Hendry and a capacitor of 5 /5 micro farate are connected in a series to a voltage source of 230 volt, 50 hz find the inpedence of the circuit


1
Expert's answer
2021-01-12T12:10:42-0500

Answer

For given LCR circuit

"X_L=2\pi fL"

"X_L=2\pi\times 50\times\frac{20}{\pi}=2000 ohm"


"X_c=\frac{1}{2\pi fc}=\frac{1}{2\pi \times50\times\frac{5}{\pi}\times10^{-6}}=2000ohm"


Impendence of circuit can be given as

"Z=\sqrt{R^2+(X_L-X_c)^2}\\Z=\sqrt{50^2+(2000-2000)^2}\\Z=50ohm"



So impendence if circuit is 50ohm